OpenStudy (anonymous):

Solve: -4(x + 1) - 3 = -3(x - 4) Answer -19 2 0 19

OpenStudy (anonymous):

-19

OpenStudy (anonymous):

Thanks (:

OpenStudy (anonymous):

-4x + -4 -3 = -3x +12 -4x - 7 = -3x + 12 x = -19

OpenStudy (anonymous):

A.-19